OpenStudy (anonymous):

simplify square root of 2800 please ASAP *************METAL ******************

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

do you know what the square root of 400 is?

OpenStudy (anonymous):

an easy way to do it is to keep taking sqrt4 out because we know it's 2, so: \[\sqrt{2800} = \sqrt{4} \times \sqrt{700} = 2\sqrt{700}\] we can do it again: \[2\sqrt{700} = 2 \times \sqrt{4} \times \sqrt{175} = 2 \times 2 \times \sqrt{175}\]

OpenStudy (anonymous):

now we have 175 which is a multiple of 25, and we know what the root of that is. i'd enjoy seeing you perform the next step :)

OpenStudy (anonymous):

it's not done though, this isn't the most simplified answer

OpenStudy (anonymous):

unless you got the point :)

OpenStudy (anonymous):

can you do it all for me please

OpenStudy (anonymous):

i guess: \[4 \sqrt{175} = 4 \times \sqrt{25} \times \sqrt{7} = 4 \times 5 \times \sqrt{7} = 20\sqrt{7}\] please return to this question in the future to understand how :)

OpenStudy (anonymous):

ok thnxx